in the households of the early settlers, porridge was a common meal. will and emma were making porridge for breakfast one morning. will contributed 200 grams of cereal toward the porridge and emma contributed 300 grams. their friend, logan joined them but he did not have any food to contribute so he gave them 50 cents for his share of food. when will and emma tried to split the 50 cents from logan they could not decide what would be fair. how do you thin they should divide the 50 cents? a fair answer is not 20 and 30 cents
Answer by Bonycells(3)

Will supplied 200 grams of cereal for porridge.
Emma supplied 300 grams of cereal for porridge.
Logan brought 50 cents but no cereal for porridge.
They all ate equal shares of the porridge.
The total amount of porridge was 500 grams of cereal.
Each person ate 500/3 grams of porridge. Each got 166-2/3 grams.
Will had 200-500/3 grams of cereal left to sell to Logan after he ate his share.
Emma had 300-500/3 grams of cereal left to sell to Logan after she ate her share.
Will had 33-1/3 grams of cereal left to sell to Logan.
Emma had 133-1/3 grams of cereal left to sell to logan.
The total cereal left over after Will and Emma ate their share was 33-1/3 plus 133-1/3 grams equalling 166-2/3 grams of cereal for Logan. All had equal shares.
Therefore, Logan must pay 50 cents for 166-2/3 grams of cereal which equals one cent for every 3-1/3 grams of cereal.
Logan must pay Will 10 cents for 33-1/3 grams of cereal.
Logan must pay Emma the balance of 40 cents for 133-1/3 grams of cereal.
That's fair.
(200p + 300p)/3 = 50c
500p/3 = 50c
50p/3 = 5c
10p/3 = 1c
3-1/3p = 1c